A Complete Bibliography of the ACM Journal on Educational Resources in Computing (JERIC)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.14

Title word cross-reference

1 [FOP06b, WDSW01].

2.0 [Fag02].

3.1 [Skr01].

8 [She02].

Abstract [SH08a]. Abstraction [SH08a]. ACM [CF01]. active [SSC01]. Ada [Fag02]. Ada/Mindstorms [Fag02]. adaptability [SFS01]. adaptive [Fis01]. Addressing [Nut06]. agent [Imb04]. aide [LB05]. Alfa [WDSW01]. Alfa-1 [WDSW01]. algorithm [Hog02, MKKN05]. Algorithmic [Arm09, Gin02]. algorithms [BA01, CH05, KPT06]. Analysis [Arm09, FM02]. Animation [Hol03, KPT06]. Animations [BGA^{+}01]. Answering [TM09]. answers [Kum05]. Approach [YKMB08, Imb04, TPQE05]. architecture [BR01a, Hog02, MAWW^{+}01, MS01, YYPA01, Yur02]. architectures [Skr01]. area [KR05]. artificial [Kum04a]. arts [TM07]. ASM [GPWD03]. assessed [MKKN05]. Assessment [JS09, BS05, BH05, DLO05, HGST05, JGB05]. assignments [BH05]. Assist [BGA^{+}01]. Assisted [JMRE02]. assurance [Bha06, FOP06a, FOP06b, GPB^{+}06]. Automated [HGST05, BH05, Kum05]. Automatic [DLO05]. automatically [MKKN05]. balancing [BC05]. Based [JMRE02, ARFILN^{+}01, BKMY03, BKMY04, DLO05, SSGS01, WJ04, MAWW^{+}01, SFS01]. Basic [Con02a]. basics [BMS02]. beginning [BMS02, GH05]. behaviors [GDB05]. binary [Gin02]. black [McK05]. board [Imb04]. boss [JGB05]. boxes [McK05].

CAN [KR05]. Case [SH08a, CR05, Hen07, Kum05]. CDM [Sut05]. Central [ZS08]. ciphers [GPG06]. circuit [Bur02]. ClassCompass [CM07]. classes [Skr01]. classroom [CR05]. classrooms [GDB05]. climates [GDB05]. closed [SSN05]. Cognitive [GKH07]. collaboration [BSWH05]. common [LK06]. Communication [GDB05]. Compiler [Gri02]. component [ARFILN^{+}01]. Computer [BGA^{+}01, DW08, GS08, GWHK05, JMRE02, SH08a, TM07, TDLG08, ZS08, BR01a, CHC03, FOP06a, FOP06b, GDB05, Har06, Hog02, Hol03, KQ04, MS01, NL03, RH06, Shu06, SPS04, Skr01, Sut05, TPQE05, WP06, WHM05, YYPA01, Yur02, FM02, Osb01]. Computer-Assisted [JMRE02]. computers [Nut06]. Computing [YKMB08, BC05, CGG^{+}05, CF01]. Concepts [ZS08, DWJD04, Hol03, MS01]. Conference [MT08a]. Constructing [GH05]. contest [VA04]. Context [TDLG08, Bis06, GKH07, TM07, YG08]. contracts [GH05]. controller [KR05]. cooperative [HKS01]. Course [Fis01, Gri02, BSWH05, Har06, Kum04a, Nut06, SKT08]. courses [Bha06]. CPU [Skr01]. Creating [LB05, GDB05]. cryptanalysis [GPG06]. Cryptography [HC02]. CS [Arm09]. CS1 [SH08b, SSN05]. CS2 [YG08]. curricula [The01]. curriculum [Con07, FOP06a, FOP06b, GPB^{+}06, McK05].

Data [SH08a, YG08]. database [Hen07]. defensive [GDB05]. Definition [SH08a]. degree [Con07]. Design [GS08, Bur02, CM07, Hen07, CR05]. Designing [Roy06]. detection [WJ04]. detective [DWJD04]. Development [GS08, CR05, SKT08]. diagramming [TPQE05]. Diagrams [JS09]. different [MKKN05]. digital [MCSF06]. disaster [LK06]. disasters [LK06]. discrete [CH05, GH05, HR05, Sut05]. Dispelling [LK06]. distance [MAWW^{+}01]. distributed [BA01, BR01b, HKS01, KPT06, SSC01]. driven [TG06]. during [SKT08].

E-Assessment [JS09]. e-learning [TG06]. early [TPQE05]. Editor [Hel01]. Editorial [CF01, FOP06a, FOP06b, HR05, Yur01]. editors [MT07a]. Education [DW08, SH08a, ZS08, ARFILN^{+}01, BC05, FM02, KQ04, Kum04b, Kum04c, MAWW^{+}01, SPS04, TPQE05, VA04]. educational [KPT06, SFS01, SSGS01, WDSW01]. educators [RH06]. Effective [Gin02]. effectiveness [GKH07]. effects [FM02]. elementary [SKT08]. embedded [RMB02]. Empirical [ZS08]. emulator [She02]. encountered [Han08]. engine [CHC03]. Engineering [Gri02, BSWH05, CR05, KQ04, VA04]. enhancing [GKH07]. environment [BKMY03, BKMY04, KPT06, Lei01, SSC01]. environments [HKS01]. evaluating [Edw03]. evaluation [BS05]. Evidence [MT07b, TM08b]. execution [BA01]. exercise [Fis01]. exercises [Bha06, BS05, MKKN05]. experience [CR05, MCSF06]. Experiences [MKKN05, HGST05]. experiential [VA04]. experiments [Bru06]. explaining [Roy06]. extensible [KPT06]. feedback [Kum05]. female [WHM05]. field [KQ04]. First [SH08b, Nut06]. First-year [SH08b]. forensics [Har06]. Foundation [ZS08]. framework [BR01b, HMHGK01, SSN05]. fully [Kum05]. future [Yur02].

Game [GS08]. gaming [GKH07]. Ganzúa [GPG06]. Gender [McK05, BC05, LB05]. gender-balancing [BC05]. Generation [Kum05]. GetSmart [MCSF06]. grade [Kum05]. graphical [Bur02]. graphics [GWHK05]. Greetings [Hel01]. Grounding [TM08a]. Guest [Hel01]. guide [Fag02].

h [MAWW^{+}01]. hands [Shu06]. hands-on [Shu06]. handy [Imb04]. HCI [YKMB08]. helps [WHM05]. High [Hog02, SH08a]. High-performance [Hog02]. hiring [CV08]. History [Hen07]. hold [Yur02]. hospital [GPB^{+}06]. host [Shu06]. Hurricane [LK06]. hypermedia [Fis01].

Implementation [SH08a]. implications [GKH07]. impressions [SH08b]. improve [SM03]. Improving [Edw03]. Individualized [BS05]. industry [CV08]. influence [LB05]. information [Bha06, Bis06, CHC03, FOP06a, FOP06b, GPB^{+}06, HKS01, Hen07]. Infrastructure [JMRE02]. Innovative [GPB^{+}06]. instruction [Bru06, GWHK05, LB05]. Instructional [DW08]. intelligence [Kum04a]. intelligent [Imb04, SSC01]. interaction [SM03]. Interactive [BA01, SFS01]. interdisciplinary [KQ04]. interpreter [CS02]. introduce [MS01]. introducing [CF01]. Introduction [Con02a, HC02, Kum04b, Kum04c]. introductory [BR01a, FM02]. IRI [MAWW^{+}01]. IRI-h [MAWW^{+}01]. ISA [DZ02]. issue [BC05, BH05, FOP06b, HR05, Kum04b, Kum04c, Yur02, FOP06b].

Java [MAWW^{+}01]. Java-based [MAWW^{+}01]. JavaScript [BMS02]. JERIC [MT07a]. Journal [CF01, MT08a].

Katrina [LK06]. kernels [RMB02]. Key [BGA^{+}01, CHC03]. knob [BR01a]. knowledge [BS05, HKS01, TG06]. knowledge-driven [TG06]. KScalar [MRL02].

L3 [Lei01]. Labels [JS09]. Laboratories [DW08, SSN05]. Laboratory [Bha06, Bru06, BMS02, VA04]. language [DH01, DWJD04, GPWD03, WJ04]. learned [Kum04a, LK06]. Learning [BGA^{+}01, JMRE02, TM08a, TDLG08, Fis01, HKS01, Lei01, MCSF06, SSC01, TG06]. Lego [GPWD03]. LEGO(R) [Imb04]. LegoSim [RMB02]. lessons [Kum04a, LK06]. level [SPS04]. liberal [TM07]. libraries [MCSF06]. library [CR05]. Linking [TM08b]. Linux [Shu06]. literate [Roy06]. Little [YYPA01]. logic [Bur02, CH05]. Logisim [Bur02]. LYDIAN [KPT06].

machine [DH01]. majors [Sut05]. Making [MT08b, GH05]. Man [YYPA01]. management [HKS01]. manual [Fag02]. MARIE [NL03]. MarieSim [NL03]. mathematics [CH05, GH05, HR05, Sut05]. may [Yur02]. means [GDB05]. mentoring [CM07]. metadata [Fis01]. microcluster [KR05]. mindstorm [GPWD03]. minimalist [TPQE05]. minimUML [TPQE05]. model [ARFILN^{+}01, Con07, GPB^{+}06, TG06]. monoalphabetic [GPG06]. Mostly [Arm09]. Moving [MCSF06]. multimedia [HMHGK01, SSGS01]. multithreaded [CMS03]. myths [LK06].

Narrating [YG08]. natural [WJ04]. network [Bru06, KR05, RH06]. networking [Hol03]. networks [Imb04]. neural [Imb04]. novice [Han08].

OAMulator [MS01]. object [Fel05, SKT08]. object-oriented [Fel05, SKT08]. online [Bha06, CH05, JGB05]. open [Lei01]. Organization [TDLG08, Skr01]. oriented [Fel05, SKT08]. other [LK06]. own [Edw03].

Pair [WHM05, Han08, SH08b]. Pair-programming [WHM05]. Parallel [KR05]. parity [LB05]. Part [TM09]. Patterns [GS08]. PDP [She02]. PDP-8 [She02]. pedagogical [CMS03]. peer [SM03]. Mindstorms [Fag02]. performance [Edw03, Hog02, MAWW^{+}01]. personalize [TG06]. perspectives [Gin02]. PIPPIN [DH01]. PL [DWJD04]. PL-detective [DWJD04]. places [Hen07]. plagiarism [WJ04]. policies [MKKN05]. polyalphabetic [GPG06]. portable [WP06, RH06]. Postroom [Osb01]. Practice [TM08a]. Preface [BH05]. present [Yur02]. Problem [JS09, Gin02]. Problems [Arm09, Han08, Kum05]. processing [DH01, KR05]. processor [WDSW01]. program [KQ04, SKT08, She02]. programmers [BMS02, GH05, Han08]. programming [BKMY03, BKMY04, BMS02, BS05, BH05, CMS03, Con02a, DWJD04, DLO05, Fel05, GPWD03, HGST05, McK05, SKT08, SH08b, WHM05]. programs [Edw03, Roy06]. Project [Gri02, CHC03, Har06]. projects [GPWD03]. promise [CGG^{+}05]. proofs [CH05]. pseudocode [Roy06]. Pthreads [RMB02]. purposes [WDSW01]. puzzles [CH05]. Pyro [BKMY03, BKMY04]. python [BKMY04, BKMY03]. python-based [BKMY04, BKMY03].

quality [Fel05]. Quantitative [Arm09, FM02]. Question [TM09]. questions [TM08b]. QuizPACK [BS05]. rationale [CR05]. Real [MT08b, YKMB08]. Realizing [CGG^{+}05]. recovery [LK06]. Reducing [SH08a]. Reduction [Arm09]. Reductive [Arm09]. reference [Fag02]. reflections [BSWH05]. relational [Hen07]. relevant [GH05]. Requirements [Con02b]. resource [MS01]. Resources [CF01, FOP06b, SFS01, SSGS01, FOP06a]. resubmission [MKKN05]. retrieval [Hen07]. Reusability [SFS01]. review [DLO05, SM03]. road [RH06]. RoboCup [SPS04]. Robot [VA04, CS02]. robotics [BKMY03, BKMY04, KQ04, Kum04b, Kum04c]. robots [FM02, Imb04, Kum04a]. role [YG08]. RTLSim [YYPA01].

Scholarship [TM08a]. School [SH08a]. Science [BGA^{+}01, SH08a, ZS08, BR01a, CHC03, GDB05, GKH07, KQ04, SPS04, Sut05, TM07, TPQE05, WHM05, FM02]. search [CHC03]. Security [DW08, Bis06, Bru06, FOP06a, FOP06b, RH06, Shu06, WP06]. see [Yur02]. SEED [DW08]. self [BS05]. self-assessment [BS05]. Sentence [WJ04]. Sentence-based [WJ04]. sequencing [Fis01]. series [BMS02]. serious [GKH07]. sharing [SSGS01]. Significant [TM09]. Sim [Skr01]. simulated [WDSW01]. simulating [Skr01]. simulation [Bur02, MKKN05, RMB02]. simulations [DH01]. simulator [BR01a, Hog02, MRL02, NL03, YYPA01]. simulators [Yur02]. Sloop [DZ02]. small [Nut06]. SMOK [DZ02]. Software [Con02b, Gri02, BSWH05, CV08, CM07, HR05]. Solutions [Arm09, RH06]. solving [Gin02]. Some [BGA^{+}01]. space [MCSF06]. Special [BC05, Yur02, BH05, HR05, Kum04b, Kum04c, FOP06b]. specialized} [Yur02]. standardized [ARFILN^{+}01]. state [SKT08]. sTeam [HKS01]. structures [YG08]. structuring [HKS01]. student [CV08, Edw03, GPWD03, MCSF06, SKT08, SM03, SH08b]. students [Edw03, WHM05]. study [Hen07, Kum05, SKT08]. submission [JGB05]. Suite [DW08]. Support [TDLG08, BR01b, HR05]. supportive [GDB05]. switch [BR01a]. system [Bur02, CM07, DWJD04, Hen07, JGB05, MAWW^{+}01]. Systems [Con02b, BR01b, CHC03, Fis01, SFS01].

Taking [RH06]. taxonomy [HMHGK01]. Teaching [Bis06, CHC03, CH05, Fel05, GS08, Gri02, Shu06, TM08a, BKMY03, BKMY04, BR01b, CR05, DWJD04, GPB^{+}06, HR05, HGST05, Imb04, MS01, YYPA01, Sut05]. team [HKS01]. team-distributed [HKS01]. term [Har06]. test [DLO05, Edw03]. test-based [DLO05]. their [Edw03]. theoretical [HMHGK01]. theory [CGG^{+}05]. ThreadMentor [CMS03]. Three [Kum04a, YYPA01]. tool [CMS03, GPG06, GPWD03, SM03, Skr01]. toolkit [DZ02]. tools [SSGS01, YYPA01]. Topics [BGA^{+}01, CHC03]. training [GKH07]. tutor [Kum05]. Types [SH08a].

UML [TPQE05]. Uncovering [CV08]. undergraduate [KQ04, Kum04b, Kum04c]. understanding [GDB05]. university [SPS04]. university-level [SPS04]. usability [CR05]. Use [SH08a, SM03]. user [Fag02]. Using [HMHGK01, SPS04, WDSW01, Fis01, Imb04, Kum04a].

values [CV08]. versatile [BKMY03, BKMY04, GPWD03]. vision [MT07a]. visual [CH05, Con02a]. visualization [CGG^{+}05]. visualizations [SKT08]. VizClass [GWHK05]. Voices [BSWH05].

Web [ARFILN^{+}01, CHC03, JMRE02, SFS01, SSGS01]. Web-based [SFS01, JMRE02, ARFILN^{+}01, SSGS01]. WebCoM [SM03]. well [Edw03]. women [BSWH05]. workshop [WP06]. World [YKMB08].

year [SH08b]. years [Kum04a].

References

Anido-Rifon:2001:SSI

[ARFILN^{+}01] L. Anido-Rifón, M. J. Fernández-Iglesias, M. Llamas-Nistal, M. Caeiro-Rodríguez, J. Santos-Gago, and J. S. Rodríguez-Estévez. A component model for standardized Web-based education. ACM Journal on Educational Resources in Computing (JERIC), 1(2es): ??, Summer 2001. CODEN ???? ISSN 1531-4278.

Armoni:2009:RCM

[Arm09] Michal Armoni. Reduction in CS: a (mostly) quantitative analysis of reductive solutions to algorithmic problems. ACM Journal on Educational Resources in Computing (JERIC), 8(4):11:1–11:??, January 2009. CODEN ???? ISSN 1531-4278.

Ben-Ari:2001:IED

[BA01] Mordechai Ben-Ari. Interactive execution of distributed algorithms. ACM Journal on Educational Resources in Computing (JERIC), 1(2es): ??, Summer 2001. CODEN ???? ISSN 1531-4278.

Bair:2005:SIG

[BC05] Bettina Bair and J. McGrath Cohoon. Special issue on gender-balancing computing education. ACM Journal on Educational Resources in Computing (JERIC), 4(1):1, March 2005. CODEN ???? ISSN 1531-4278.

Balci:2001:AAL

[BGA^{+}01] Osman Balci, William S.

Gilley, Robin J. Adams, Emre Tunar, and N. Dwight Barnette. Animations to assist learning some key computer science topics. ACM Journal on Educational Resources in Computing (JERIC), 1(2es): ??, Summer 2001. CODEN ???? ISSN 1531-4278.

Brusilovsky:2005:PSI

[BH05] Peter Brusilovsky and Colin Higgins. Preface to the special issue on automated assessment of programming assignments. ACM Journal on Educational Resources in Computing (JERIC), 5(3):??, September 2005. CODEN ???? ISSN 1531-4278.

Bhagyavati:2006:LEO

[Bha06] Bhagyavati. Laboratory exercises in online information assurance courses. ACM Journal on Educational Resources in Computing (JERIC), 6(4): 4:1–4:??, December 2006. CO-DEN ???? ISSN 1531-4278.

Bishop:2006:TCI

[Bis06] Matt Bishop. Teaching context in information security. ACM Journal on Educational Resources in Computing (JERIC), 6(3):3:1–3:??, September 2006. CODEN ???? ISSN 1531-4278.

Blank:2003:PPB

[BKMY03] Douglas Blank, Deepak Kumar, Lisa Meeden, and Holly Yanco. Pyro: a Python-based versatile programming environment for teaching robotics. ACM Journal on Educational Resources in Computing (JERIC), 3(4):1–15, December 2003. CODEN ???? ISSN 1531-4278.

Blank:2004:PPB

[BKMY04] Douglas Blank, Deepak Kumar, Lisa Meeden, and Holly Yanco. Pyro: a python-based versatile programming environment for teaching robotics. ACM Journal on Educational Resources in Computing (JERIC), 4(3):1–15, September 2004. CODEN ???? ISSN 1531-4278.

Brady:2002:JPB

[BMS02] Alyce Brady, R. C. McDowell, and Kelly Schultz. JavaScript programming basics: a laboratory series for beginning programmers. ACM Journal on Educational Resources in Computing (JERIC), 2(2):1, June 2002. CODEN ???? ISSN 1531-4278.

Braught:2001:KSC

[BR01a] Grant Braught and David Reed. The knob & switch computer: a computer architecture simulator for introductory computer science. ACM Journal on Educational Resources in Computing (JERIC), 1(4): 31–45, December 2001. CO-DEN ???? ISSN 1531-4278.

Burger:2001:FST

[BR01b] C. Burger and K. Rothermel. A framework to support teaching in distributed systems. ACM Journal on Educational Resources in Computing (JERIC), 1(1es):??, March 2001. CODEN ???? ISSN 1531-4278.

Brustoloni:2006:LEN

[Bru06] José Carlos Brustoloni. Laboratory experiments for network security instruction. ACM Journal on Educational Resources in Computing (JERIC), 6(4):5:1–5:??, December 2006. CODEN ???? ISSN 1531-4278.

Brusilovsky:2005:IES

[BS05] Peter Brusilovsky and Sergey Sosnovsky. Individualized exercises for self-assessment of programming knowledge: An evaluation of QuizPACK. ACM Journal on Educational Resources in Computing (JERIC), 5(3):??, September 2005. CODEN ???? ISSN 1531-4278.

Berenson:2005:VWS

[BSWH05] Sarah B. Berenson, Kelli M.

Slaten, Laurie Williams, and Chih-Wei Ho. Voices of women in a software engineering course: reflections on collaboration. ACM Journal on Educational Resources in Computing (JERIC), 4(1):1– 18, March 2005. CODEN ???? ISSN 1531-4278.

Burch:2002:LGS

[Bur02]  Carl Burch.  Logisim: a

graphical system for logic circuit design and simulation. ACM Journal on Educational Resources in Computing (JERIC), 2(1):5–16, March 2002. CODEN ???? ISSN 1531-4278.

Cassel:2001:EIA

[CF01] Lillian N. Cassel and Edward A. Fox. Editorial: introducing the ACM Journal on Resources in Computing. ACM Journal on Educational Resources in Computing (JERIC), 1(1es):??, March 2001. CODEN ???? ISSN 1531-4278.

Cogliati:2005:RPV

[CGG^{+}05] Joshua J. Cogliati, Frances W. Goosey, Michael T. Grinder, Bradley A. Pascoe, Rockford J. ROSS, and Cheston J. Williams. Realizing the promise of visualization in the theory of computing. ACM Journal on Educational Resources in Computing (JERIC), 5(2):1–17, June 2005. CODEN ???? ISSN 1531-4278.

Cigas:2005:TPA

[CH05] John Cigas and Wen-Jung Hsin. Teaching proofs and algorithms in discrete mathematics with online visual logic puzzles. ACM Journal on Educational Resources in Computing (JERIC), 5(2):1–12, June 2005. CODEN ???? ISSN 1531-4278.

Chau:2003:TKT

[CHC03] Michael Chau, Zan Huang, and Hsinchun Chen. Teaching key topics in computer science and information systems through a Web search engine project. ACM Journal on Educational Resources in Computing (JERIC), 3(3): 1–14, September 2003. CO-DEN ???? ISSN 1531-4278.

Coelho:2007:CSD

[CM07] Wesley Coelho and Gail Murphy. ClassCompass: a software design mentoring system. ACM Journal on Educational Resources in Computing (JERIC), 7(1):2:1–2:??, March 2007. CODEN ???? ISSN 1531-4278.

Carr:2003:TPT

[CMS03] Steve Carr, Jean Mayo, and Ching-Kuang Shene. Thread-Mentor: a pedagogical tool for multithreaded programming. ACM Journal on Educational Resources in Computing (JERIC), 3(1):1–30, March 2003. CODEN ???? ISSN 1531-4278.

Conn:2002:IPV

[Con02a] Richard Conn. Introduction to programming with Visual Basic. ACM Journal on Educational Resources in Computing (JERIC), 2(3):1, September 2002. CODEN ???? ISSN 1531-4278.

Conn:2002:SSR

[Con02b] Richard Conn. Software systems requirements. ACM Journal on Educational Resources in Computing (JERIC), 2(4):1, December 2002. CODEN ???? ISSN 1531-4278.

Consortium:2007:MCL

[Con07] Liberal Arts Computer Science Consortium. A 2007 model curriculum for a liberal arts degree in computer science. ACM Journal on Educational Resources in Computing (JERIC), 7(2):2:1–2:??, June 2007. CODEN ???? ISSN 1531-4278.

Carroll:2005:CLT

[CR05] John M. Carroll and Mary Beth Rosson. A case library for teaching usability engineering: Design rationale, development, and classroom experience. ACM Journal on Educational Resources in Computing (JERIC), 5(1):1–22, March 2005. CODEN ???? ISSN 1531-4278.

Carlisle:2002:RI

[CS02] Martin C. Carlisle and Keith Shomper. ROBOT interpreter. ACM Journal on Educational Resources in Computing (JERIC), 2(2):2, June 2002. CODEN ???? ISSN 1531-4278.

Chinn:2008:USV

[CV08] Donald Chinn and Tammy Vandegrift. Uncovering student values for hiring in the software industry. ACM Journal on Educational Resources in Computing (JERIC), 7(4): 4:1–4:??, January 2008. CO-DEN ???? ISSN 1531-4278.

Decker:2001:PMS

[DH01] Rick Decker and Stuart Hirshfield. The PIPPIN machine: simulations of language processing. ACM Journal on Educational Resources in Computing (JERIC), 1(4):4–17, December 2001. CODEN ???? ISSN 1531-4278.

Douce:2005:ATB

[DLO05] Christopher Douce, David Livingstone, and James Orwell. Automatic test-based assessment of programming: a review. ACM Journal on Educational Resources in Computing (JERIC), 5(3):??, September 2005. CODEN ???? ISSN 1531-4278.

Du:2008:SSI

[DW08] Wenliang Du and Ronghua Wang. SEED: a suite of instructional laboratories for computer security education. ACM Journal on Educational Resources in Computing (JERIC), 8(1):3:1–3:??, March 2008. CODEN ???? ISSN 1531-4278.

Diwan:2004:PDS

[DWJD04] Amer Diwan, William M. Waite, Michele H. Jackson, and Jacob Dickerson. PLdetective: a system for teaching programming language concepts. ACM Journal on Educational Resources in Computing (JERIC), 4(4):1–22, December 2004. CODEN ???? ISSN 1531-4278.

Dugan:2002:SIS

[DZ02] B. Dugan and J. Zahorjan. The Sloop ISA and the SMOK toolkit. ACM Journal on Educational Resources in Computing (JERIC), 2(1):49–71, March 2002. CODEN ???? ISSN 1531-4278.

Edwards:2003:ISP

[Edw03] Stephen H. Edwards. Improving student performance by evaluating how well students test their own programs. ACM Journal on Educational Resources in Computing (JERIC), 3(3):1–24, September 2003. CODEN ???? ISSN 1531-4278.

Fagin:2002:AMU

[Fag02] Barry Fagin. Ada/Mindstorms 2.0 user’s guide and reference manual. ACM Journal on Educational Resources in Computing (JERIC), 2(2):3, June 2002. CODEN ???? ISSN 1531-4278.

Feldman:2005:TQO

[Fel05] Yishai A. Feldman. Teaching quality object-oriented programming. ACM Journal on Educational Resources in Computing (JERIC), 5(1):1– 16, March 2005. CODEN ???? ISSN 1531-4278.

Fischer:2001:CES

[Fis01]  Stephan Fischer.  Course

and exercise sequencing using metadata in adaptive hypermedia learning systems. ACM Journal on Educational Resources in Computing (JERIC), 1(1es):??, March 2001. CODEN ???? ISSN 1531-4278.

Fagin:2002:QAE

[FM02] Barry S. Fagin and Laurence Merkle. Quantitative analysis of the effects of robots on introductory Computer Science education. ACM Journal on Educational Resources in Computing (JERIC), 2(4):1–17, December 2002. CODEN ???? ISSN 1531-4278.

Frincke:2006:ERC

[FOP06a] D. Frincke, S. Oudekirk, and B. Popovsky. Editorial: Resources for the computer security and information assurance curriculum. ACM Journal on Educational Resources in Computing (JERIC), 6(4): 1:1–1:??, December 2006. CO-DEN ???? ISSN 1531-4278.

Frincke:2006:ESI

[FOP06b] D. Frincke, S. Oudekirk, and B. Popovsky. Editorial: Special issue on resources for the computer security and information assurance curriculum: Issue 1. ACM Journal on Educational Resources in Computing (JERIC), 6(3):1:1–1:??, September 2006. CODEN ???? ISSN 1531-4278.

Garvin-Doxas:2005:CCS

[GDB05] Kathy Garvin-Doxas and Lecia J. Barker. Communication in computer science classrooms: understanding defensive climates as a means of creating supportive behaviors. ACM Journal on Educational Resources in Computing (JERIC), 4(1):1–18, March 2005. CODEN ???? ISSN 1531-4278.

Gegg-Harrison:2005:CCM

[GH05] Timothy S. Gegg-Harrison. Constructing contracts: Making discrete mathematics relevant to beginning programmers. ACM Journal on Educational Resources in Computing (JERIC), 5(2):1–28, June 2005. CODEN ???? ISSN 1531-4278.

Ginat:2002:EBP

[Gin02] David Ginat. Effective binary perspectives in algorithmic problem solving. ACM Journal on Educational Resources in Computing (JERIC), 2(2): 4, June 2002. CODEN ???? ISSN 1531-4278.

Greitzer:2007:CSI

[GKH07] Frank L. Greitzer, Olga Anna Kuchar, and Kristy Huston.

Cognitive science implications for enhancing training effectiveness in a serious gaming context. ACM Journal on Educational Resources in Computing (JERIC), 7(3):2:1–2:??, November 2007. CODEN ???? ISSN 1531-4278.

Goel:2006:IMI

[GPB^{+}06] Sanjay Goel, Damira Pon, Peter Bloniarz, Robert Bangert-Drowns, George Berg, Vince Delio, Laura Iwan, Thomas Hurbanek, Sandoor P. Schuman, Jagdish Gangolly, Adnan Baykal, and Jon Hobbs. Innovative model for information assurance curriculum: A teaching hospital. ACM Journal on Educational Resources in Computing (JERIC), 6(3): 2:1–2:??, September 2006. CO-DEN ???? ISSN 1531-4278.

Garcia-Pasquel:2006:GCT

[GPG06] Jesús Adolfo García-Pasquel and José Galaviz. Ganzúa: a cryptanalysis tool for monoalphabetic and polyalphabetic ciphers. ACM Journal on Educational Resources in Computing (JERIC), 6(3):4:1–4:??, September 2006. CODEN ???? ISSN 1531-4278.

Gilder:2003:VTS

[GPWD03] Jason Gilder, Michael Peterson, Jason Wright, and Travis Doom. A versatile tool for student projects: an ASM programming language for the Lego mindstorm. ACM Journal on Educational Resources in Computing (JERIC), 3(1): 1–14, March 2003. CODEN ???? ISSN 1531-4278.

Griswold:2002:TSE

[Gri02] William G. Griswold. Teaching software engineering in a compiler project course. ACM Journal on Educational Resources in Computing (JERIC), 2(4):1–18, December 2002. CODEN ???? ISSN 1531-4278.

Gestwicki:2008:TDP

[GS08] Paul Gestwicki and Fu-Shing Sun. Teaching design patterns through computer game development. ACM Journal on Educational Resources in Computing (JERIC), 8(1):2:1–2:??, March 2008. CODEN ???? ISSN 1531-4278.

Grimes:2005:CGI

[GWHK05] Douglas Grimes, Mark Warschauer, Tara Hutchinson, and Falko Kuester. Computer graphics instruction in VizClass. ACM Journal on Educational Resources in Computing (JERIC), 5(4):3:1–3:??, December 2005. CODEN ???? ISSN 1531-4278.

Hanks:2008:PEN

[Han08] Brian Hanks. Problems encountered by novice pair programmers. ACM Journal on Educational Resources in Computing (JERIC), 7(4):2:1– 2:??, January 2008. CODEN ???? ISSN 1531-4278.

Harrison:2006:TPC

[Har06] Warren Harrison. A term project for a course on computer forensics. ACM Journal on Educational Resources in Computing (JERIC), 6(3): 6:1–6:??, September 2006. CO-DEN ???? ISSN 1531-4278.

Humphries:2002:IC

[HC02] Jeffrey W. Humphries and Martin C. Carlisle. Introduction to Cryptography. ACM Journal on Educational Resources in Computing (JERIC), 2(3):2, September 2002. CODEN ???? ISSN 1531-4278.

Heller:2001:GGE

[Hel01] R. S. Heller. Greetings from the Guest Editor. ACM Journal on Educational Resources in Computing (JERIC), 1(1es): ??, March 2001. CODEN ???? ISSN 1531-4278.

Hendry:2007:HPC

[Hen07] David G. Hendry. History places: a case study for relational database and information retrieval system design. ACM Journal on Educational Resources in Computing (JERIC), 7(1):3:1–3:??, March 2007. CODEN ???? ISSN 1531-4278.

Higgins:2005:AAE

[HGST05] Colin A. Higgins, Geoffrey Gray, Pavlos Symeonidis, and Athanasios Tsintsifas. Automated assessment and experiences of teaching programming. ACM Journal on Educational Resources in Computing (JERIC), 5(3):??, September 2005. CODEN ???? ISSN 1531-4278.

Hampel:2001:SSI

[HKS01] Thorsten Hampel and Reinhard Keil-Slawik. sTeam: structuring information in team-distributed knowledge management in cooperative learning environments. ACM Journal on Educational Resources in Computing (JERIC), 1(2es):??, Summer 2001. CO-DEN ???? ISSN 1531-4278.

Heller:2001:UTM

[HMHGK01] Rachelle S. Heller, C. Dianne Martin, Nuzi Haneef, and Sonja Gievska-Krliu. Using a theoretical multimedia taxonomy framework. ACM Journal on Educational Resources in Computing (JERIC), 1(1es): ??, March 2001. CODEN ???? ISSN 1531-4278.

Hoganson:2002:HPC

[Hog02] Kenneth E. Hoganson. Highperformance computer architecture and algorithm simulator. ACM Journal on Educational Resources in Computing (JERIC), 2(1):131–148, March 2002. CODEN ???? ISSN 1531-4278.

Holliday:2003:ACN

[Hol03] Mark A. Holliday. Animation of computer networking concepts. ACM Journal on Educational Resources in Computing (JERIC), 3(2):1–26, June 2003. CODEN ???? ISSN 1531-4278.

Harvey:2005:ESI

[HR05] Valerie J. Harvey and Susan H.

Rodger. Editorial for the special issue on software support for teaching discrete mathematics. ACM Journal on Educational Resources in Computing (JERIC), 5(2):1–16, June 2005. CODEN ???? ISSN 1531-4278.

Imberman:2004:IAA

[Imb04] Susan P. Imberman. An intelligent agent approach for teaching neural networks using LEGO(R) handy board robots. ACM Journal on Educational Resources in Computing (JERIC), 4(3):1–12, September 2004. CODEN ???? ISSN 1531-4278.

Joy:2005:BOS

[JGB05] Mike Joy, Nathan Griffiths, and Russell Boyatt. The boss online submission and assessment system. ACM Journal on Educational Resources in Computing (JERIC), 5(3):??, September 2005. CODEN ???? ISSN 1531-4278.

Joy:2002:IWB

[JMRE02] Mike Joy, Boris Muzykantskii, Simon Rawles, and Michael Evans. An infrastructure for Web-based computer-assisted learning. ACM Journal on Educational Resources in Computing (JERIC), 2(4):1–19, December 2002. CODEN ???? ISSN 1531-4278.

Jayal:2009:PLA

[JS09] Ambikesh Jayal and Martin Shepperd. The problem of labels in E-assessment of diagrams. ACM Journal on Educational Resources in Computing (JERIC), 8(4):12:1–12:??, January 2009. CODEN ???? ISSN 1531-4278.

Koldehofe:2006:LEE

[KPT06] Boris Koldehofe, Marina Papatriantafilou, and Philippas Tsigas. LYDIAN: An extensible educational animation environment for distributed algorithms. ACM Journal on Educational Resources in Computing (JERIC), 6(2):1:1–1:??, June 2006. CODEN ???? ISSN 1531-4278.

Kitts:2004:IFR

[KQ04] Christopher Kitts and Neil Quinn. An interdisciplinary field robotics program for undergraduate computer science and engineering education. ACM Journal on Educational Resources in Computing (JERIC), 4(2):3, June 2004. CODEN ???? ISSN 1531-4278.

Kuban:2005:CMP

[KR05] Paul A. Kuban and Rammohan K. Ragade. The CAN microcluster: Parallel processing over the controller area network. ACM Journal on Educational Resources in Computing (JERIC), 5(1):1–12, March 2005. CODEN ???? ISSN 1531-4278.

Kumar:2004:TYU

[Kum04a] Amruth N. Kumar. Three years of using robots in an artificial intelligence course: lessons learned. ACM Journal on Educational Resources in Computing (JERIC), 4(3): 1–15, September 2004. CO-DEN ???? ISSN 1531-4278.

Kumar:2004:ISIa

[Kum04b] Deepak Kumar. Introduction to special issue on robotics in undergraduate education. ACM Journal on Educational Resources in Computing (JERIC), 4(2):1, June 2004. CODEN ???? ISSN 1531-4278.

Kumar:2004:ISIb

[Kum04c] Deepak Kumar. Introduction to special issue on robotics in undergraduate education. ACM Journal on Educational Resources in Computing (JERIC), 4(3):1, September 2004. CODEN ???? ISSN 1531-4278.

Kumar:2005:GPA

[Kum05] Amruth N. Kumar. Generation of problems, answers, grade, and feedback—case study of a fully automated tutor. ACM Journal on Educational Resources in Computing (JERIC), 5(3):??, September 2005. CODEN ???? ISSN 1531-4278.

Lester:2005:CGP

[LB05] Cynthia Y. Lester and Marcus Brown. Creating gender parity: an instruction aide’s influence. ACM Journal on Educational Resources in Computing (JERIC), 4(1):1–14, March 2005. CODEN ???? ISSN 1531-4278.

Leidig:2001:LTO

[Lei01] Torsten Leidig. L3—towards an open learning environment. ACM Journal on Educational Resources in Computing (JERIC), 1(1es):??, March 2001. CODEN ???? ISSN 1531-4278.

Landry:2006:DCD

[LK06] Brett J. L. Landry and M. Scott Koger. Dispelling 10 common disaster recovery myths: Lessons learned from Hurricane Katrina and other disasters. ACM Journal on Educational Resources in Computing (JERIC), 6(4):6:1–6:??, December 2006. CODEN ???? ISSN 1531-4278.

Maly:2001:IHJ

[MAWW^{+}01] R. Maly, H. Abdel-Wahab, C. Wild, C. M. Overstreet, A. Gupta, A. Abdel-Hamid, S. Ghanem, A. Gonzalez, and X. Zhu. IRI-h, a Javabased distance education system: architecture and performance. ACM Journal on Educational Resources in Computing (JERIC), 1(1es):??, March 2001. CODEN ???? ISSN 1531-4278.

McKenna:2005:GBB

[McK05] Peter McKenna. Gender and black boxes in the programming curriculum. ACM Journal on Educational Resources in Computing (JERIC), 4(1): 1–12, March 2005. CODEN ???? ISSN 1531-4278.

Marshall:2006:MDL

[MCSF06] Byron B. Marshall, Hsinchun Chen, Rao Shen, and Edward A. Fox. Moving digital libraries into the student learning space: The GetSmart experience. ACM Journal on Educational Resources in Computing (JERIC), 6(1):2:1–2:??, March 2006. CODEN ???? ISSN 1531-4278.

Malmi:2005:EAA

[MKKN05] Lauri Malmi, Ville Karavirta, Ari Korhonen, and Jussi Nikander. Experiences on automatically assessed algorithm simulation exercises with different resubmission policies. ACM Journal on Educational Resources in Computing (JERIC), 5(3):??, September 2005. CODEN ???? ISSN 1531-4278.

Moure:2002:KS

[MRL02] J. C. Moure, Dolores I. Rexachs, and Emilio Luque. The KScalar simulator. ACM Journal on Educational Resources in Computing (JERIC), 2(1):

73–116, March 2002. CODEN ???? ISSN 1531-4278.

Menczer:2001:OTR

[MS01] Filippo Menczer and Alberto Maria Segre. OAMulator: a teaching resource to introduce computer architecture concepts. ACM Journal on Educational Resources in Computing (JERIC), 1(4):18– 30, December 2001. CODEN ???? ISSN 1531-4278.

McCartney:2007:ENV

[MT07a] Robert McCartney and Josh Tenenberg. From the editors: a new vision for JERIC. ACM Journal on Educational Resources in Computing (JERIC), 7(1):1:1–1:??, March 2007. CODEN ???? ISSN 1531-4278.

McCartney:2007:WE

[MT07b] Robert McCartney and Josh Tenenberg. Why evidence? ACM Journal on Educational Resources in Computing (JERIC), 7(3):1:1–1:??, November 2007. CODEN ???? ISSN 1531-4278.

McCartney:2008:CJ

[MT08a] Robert McCartney and Josh Tenenberg. From conference to journal. ACM Journal on Educational Resources in Computing (JERIC), 8(1):1:1–1:??, March 2008. CODEN ???? ISSN 1531-4278.

McCartney:2008:MIR

[MT08b] Robert McCartney and Josh Tenenberg. Making it real.

ACM Journal on Educational Resources in Computing (JERIC), 8(3):7:1–7:??, October 2008. CODEN ???? ISSN 1531-4278.

Null:2003:MMC

[NL03] Linda Null and Julia Lobur.

MarieSim: The MARIE computer simulator. ACM Journal on Educational Resources in Computing (JERIC), 3(2): 1–29, June 2003. CODEN ???? ISSN 1531-4278.

Nutt:2006:ASC

[Nut06] Gary Nutt. Addressing small computers in the first OS course. ACM Journal on Educational Resources in Computing (JERIC), 6(2):2:1–2:??, June 2006. CODEN ???? ISSN 1531-4278.

Osborne:2001:PC

[Osb01] Hugh Osborne. The Postroom Computer. ACM Journal on Educational Resources in Computing (JERIC), 1(4):81– 110, December 2001. CODEN ???? ISSN 1531-4278.

Rosenberg:2006:TNR

[RH06] Timothy Rosenberg and Lance J.

Hoffman. Taking the network on the road: Portable network solutions for computer security educators. ACM Journal on Educational Resources in Computing (JERIC), 6(4):2:1– 2:??, December 2006. CODEN ???? ISSN 1531-4278.

Roblitz:2002:LSE

[RMB02] Thomas Röblitz, Frank Mueller, and Oliver Bühn. LegoSim: simulation of embedded kernels over Pthreads. ACM Journal on Educational Resources in Computing (JERIC), 2(1): 117–130, March 2002. CODEN ???? ISSN 1531-4278.

Roy:2006:DEP

[Roy06] Geoffrey G. Roy. Designing and explaining programs with a literate pseudocode. ACM Journal on Educational Resources in Computing (JERIC), 6(1):1:1–1:??, March 2006. CODEN ???? ISSN 1531-4278.

Saddik:2001:RAI

[SFS01] Abdulmotaleb El Saddik, Stephan Fischer, and Ralf Steinmetz. Reusability and adaptability of interactive resources in Web-based educational systems. ACM Journal on Educational Resources in Computing (JERIC), 1(1es): ??, March 2001. CODEN ???? ISSN 1531-4278.

Sakhnini:2008:RAH

[SH08a] Victoria Sakhnini and Orit Hazzan. Reducing abstraction in high school computer science education: The case of definition, implementation, and use of abstract data types. ACM Journal on Educational Resources in Computing (JERIC), 8(2):5:1–5:??, June 2008. CODEN ???? ISSN 1531-4278.

Simon:2008:FYS

[SH08b] Beth Simon and Brian Hanks.

First-year students’ impressions of pair programming in CS1. ACM Journal on Educational Resources in Computing (JERIC), 7(4):5:1–5:??, January 2008. CODEN ???? ISSN 1531-4278.

Shelburne:2002:PEP

[She02] Brian J. Shelburne. A PDP-8 emulator program. ACM Journal on Educational Resources in Computing (JERIC), 2(1): 17–47, March 2002. CODEN ???? ISSN 1531-4278.

Shumba:2006:THL

[Shu06] Rose Shumba. Teaching handson Linux host computer security. ACM Journal on Educational Resources in Computing (JERIC), 6(3):5:1–5:??, September 2006. CODEN ???? ISSN 1531-4278.

Skrien:2001:CST

[Skr01] Dale Skrien. CPU Sim 3.1:

a tool for simulating computer architectures for computer organization classes. ACM Journal on Educational Resources in Computing (JERIC), 1(4): 46–59, December 2001. CO-DEN ???? ISSN 1531-4278.

URL http://www.cs.colby.
edu/djskrien/CPUSim.

Sajaniemi:2008:SDS

[SKT08] Jorma Sajaniemi, Marja Kuittinen, and Taina Tikansalo. A study of the development of students’ visualizations of program state during an elementary object-oriented programming course. ACM Journal on Educational Resources in Computing (JERIC), 7(4):3:1– 3:??, January 2008. CODEN ???? ISSN 1531-4278.

Silva:2003:WTU

[SM03] Elaine Silva and Dilvan Moriera. WebCoM: a tool to use peer review to improve student interaction. ACM Journal on Educational Resources in Computing (JERIC), 3(1): 1–14, March 2003. CODEN ???? ISSN 1531-4278.

Sklar:2004:URU

[SPS04] Elizabeth Sklar, Simon Parsons, and Peter Stone. Using RoboCup in universitylevel computer science education. ACM Journal on Educational Resources in Computing (JERIC), 4(2):4, June 2004. CODEN ???? ISSN 1531-4278.

Shang:2001:IDE

[SSC01] Yi Shang, Hongchi Shi, and Su-Shing Chen. An intelligent distributed environment for active learning. ACM Journal on Educational Resources in Computing (JERIC), 1(2es): ??, Summer 2001. CODEN ???? ISSN 1531-4278.

Shirmohammadi:2001:WBM

[SSGS01] Shervin Shirmohammadi, Abdulmotaleb El Saddik, Nicolas D. Georganas, and Ralf Steinmetz. Web-based multimedia tools for sharing educational resources. ACM Journal on Educational Resources in Computing (JERIC), 1(1es): ??, March 2001. CODEN ???? ISSN 1531-4278.

Soh:2005:FCC

[SSN05] Leen-Kiat Soh, Ashok Samal, and Gwen Nugent. A framework for CS1 closed laboratories. ACM Journal on Educational Resources in Computing (JERIC), 5(4):2:1–2:??, December 2005. CODEN ???? ISSN 1531-4278.

Sutner:2005:CTD

[Sut05] Klaus Sutner. CDM: Teaching discrete mathematics to computer science majors. ACM Journal on Educational Resources in Computing (JERIC), 5(2):1–11, June 2005. CODEN ???? ISSN 1531-4278.

Tew:2008:CSL

[TDLG08] Allison Elliott Tew, Brian Dorn, William D. Leahy, Jr., and Mark Guzdial. Context as support for learning computer organization. ACM Journal on Educational Resources in Computing (JERIC), 8(3):8:1– 8:??, October 2008. CODEN ???? ISSN 1531-4278.

Teo:2006:KDM

[TG06] Chao Boon Teo and Robert Kheng Leng Gay. A knowledgedriven model to personalize elearning. ACM Journal on Educational Resources in Computing (JERIC), 6(1):3:1–3:??, March 2006. CODEN ???? ISSN 1531-4278.

JTFCC:2001:CC

[The01] The Joint Task Force on Computing Curricula. Computing curricula 2001. ACM Journal on Educational Resources in Computing (JERIC), 1(3es): ??, Fall 2001. CODEN ???? ISSN 1531-4278.

Tenenberg:2007:CSL

[TM07] Josh Tenenberg and Robert McCartney. Computer science in a liberal arts context. ACM Journal on Educational Resources in Computing (JERIC), 7(2):1:1–1:??, June 2007. CODEN ???? ISSN 1531-4278.

Tenenberg:2008:GST

[TM08a] Josh Tenenberg and Robert McCartney. Grounding the scholarship of teaching and learning in practice. ACM Journal on Educational Resources in Computing (JERIC), 8(2):4:1–4:??, June 2008. CO-DEN ???? ISSN 1531-4278.

Tenenberg:2008:LQE

[TM08b] Josh Tenenberg and Robert McCartney. Linking questions and evidence. ACM Journal on Educational Resources in Computing (JERIC), 7(4):1:1– 1:??, January 2008. CODEN ???? ISSN 1531-4278.

Tenenberg:2009:APS

[TM09] Josh Tenenberg and Robert McCartney. Answering part of the significant question. ACM Journal on Educational Resources in Computing (JERIC), 8(4):10:1–10:??, January 2009. CODEN ???? ISSN 1531-4278.

Turner:2005:MMA

[TPQE05] Scott A. Turner, Manuel A.

Pérez-Quiñones, and Stephen H. Edwards. minimUML: a minimalist approach to UML diagramming for early computer science education. ACM Journal on Educational Resources in Computing (JERIC), 5(4): 1:1–1:??, December 2005. CO-DEN ???? ISSN 1531-4278.

Verner:2004:RCL

[VA04] Igor M. Verner and David J.

Ahlgren. Robot contest as a laboratory for experiential engineering education. ACM Journal on Educational Resources in Computing (JERIC), 4(2):2, June 2004. CODEN ???? ISSN 1531-4278.

Wainer:2001:UAS

[WDSW01] Gabriel A. Wainer, Sergio Daicz, Luis F. De Simoni, and Demian Wassermann. Using the Alfa-1 simulated processor for educational purposes. ACM Journal on Educational Resources in Computing (JERIC), 1(4):111–151, December 2001. CODEN ???? ISSN 1531-4278.

Werner:2005:PPH

[WHM05] Linda L. Werner, Brian Hanks, and Charlie McDowell. Pair-programming helps female computer science students. ACM Journal on Educational Resources in Computing (JERIC), 4(1):1–8, March 2005. CODEN ???? ISSN 1531-4278.

White:2004:SBN

[WJ04] Daniel R. White and Mike S. Joy. Sentence-based natural language plagiarism detection. ACM Journal on Educational Resources in Computing (JERIC), 4(4):1–20, December 2004. CODEN ???? ISSN 1531-4278.

Wagner:2006:PCS

[WP06] Paul J. Wagner and Andrew T. Phillips. A portable computer security workshop. ACM Journal on Educational Resources in Computing (JERIC), 6(4): 3:1–3:??, December 2006. CO-DEN ???? ISSN 1531-4278.

Yarosh:2008:NDS

[YG08] Svetlana Yarosh and Mark Guzdial. Narrating data structures: The role of context in CS2. ACM Journal on Educational Resources in Computing (JERIC), 7(4):6:1–6:??, January 2008. CODEN ???? ISSN 1531-4278.

Yardi:2008:HAC

[YKMB08] Sarita Yardi, Pamela Krolikowski, Taneshia Marshall, and Amy Bruckman. An HCI approach to computing in the real world. ACM Journal on Educational Resources in Computing (JERIC), 8(3):9:1– 9:??, October 2008. CODEN ???? ISSN 1531-4278.

Yurcik:2001:E

[Yur01]  William Yurcik.  Edito-

rial. ACM Journal on Educational Resources in Computing (JERIC), 1(4):1–3, December 2001. CODEN ???? ISSN 1531-4278.

Yurcik:2002:SIS

[Yur02] William (Bill) Yurcik. Special issue on specialized computer architecture simulators that see the present and may hold the future. ACM Journal on Educational Resources in Computing (JERIC), 2(1): 1–3, March 2002. CODEN ???? ISSN 1531-4278.

Yehezkel:2001:TST

[YYPA01] Cecile Yehezkel, William Yurcik, Murray Pearson, and Dean Armstrong. Three simulator tools for teaching computer architecture: Little Man computer, and RTL-Sim. ACM Journal on Educational Resources in Computing (JERIC), 1(4):60–80, December 2001. CODEN ???? ISSN 1531-4278.

Zendler:2008:EFC

[ZS08] Andreas Zendler and Christian Spannagel. Empirical foundation of central concepts for computer science education. ACM Journal on Educational Resources in Computing (JERIC), 8(2):6:1–6:??, June 2008. CODEN ???? ISSN 1531-4278.